OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1? I know the answer is (-1,-1), but I don't know how.

OpenStudy (luigi0210):

Yea, you could of it many ways. But use a triangle.

OpenStudy (anonymous):

A triangle?

OpenStudy (anonymous):

this is an equilateral triangle, so the y value is the same as the x value not equilateral, "isosceles" angles are π4,π2,π4 so two legs are equal

OpenStudy (anonymous):

and welcome 2 os

OpenStudy (luigi0210):

Yup, we know \[\LARGE \theta=-\frac{3\pi}{4}\]  Since we know that the one side is -1, we just need to find x. Also, anything that's a radian with 4 in it is 45 degrees, and a 45 degree triangle:  So they would be the same. Plus only tan/cot is positive in the third quad.